8. Prove that max { x ( 2 -*): x is a real number} < 1 using the MAX/MIN method. 9 Sunoco thot Sand Taro qubooto ful
Answer #1

y = x (2-1) 21 ] y= 20-12 y = 2-21 for max or min glo 2-24 0 In=1 yll = – 2 to at naj so local maxima Imax (n=1 = 1(2-1) Imaplease upvote